# Relative topological principality and the ideal intersection property for groupoid C*-algebras

### v1

@@ Abstract.
${}^{1}$ Supported by NSERC Discovery Grant RGPIN-2021-02459
${}^{2}$ Supported by NSERC Discovery Grant RGPIN-2017-04052
@@ 1. Introduction
One often wants a method for determining the injectivity of ***-homorphisms with domain a fixed C*-algebra $A$. In some situations there may be a C*-subalgebra $B$ of $A$ such that any ***-homomorphism is injective on $A$ if and only if it is injective on $B$; this is particularly useful if the subalgebra $B$ is more easily understood than the whole algebra $A$. Equivalently, we say that a C*-subalgebra $B$ of $A$ detects ideals in $A$, or has the ideal intersection property in $A$, if the only (closed 2-sided) ideal $J$ of $A$ such that $B\cap J=\{0\}$ is the zero ideal. If $B$ detects ideals in $A$, then we can verify injectivity of a ***-homomorphism on $A$ by checking that it is injective on $B$. There are many cases known where certain types of C*-algebras have important subalgebras with the ideal intersection property, e.g. [KawamuraTomiyama, ArchboldSpielberg, ExelLacaQuigg, exelNonHausdorffEtaleGroupoids2011, LS, brown2015cartan, Starling, kennedyIdealIntersectionProperty2021] (see Section 3 below).
In this paper we consider the ideal intersection property for subalgebras and, more generally, families of subalgebras of $C^{*}_{r}(G)$, where $G$ is a second-countable Hausdorff étale groupoid. We introduce the notion of a groupoid being relatively topologically principal with respect to a family of open subgroupoids. Our main result is that if $G$ is relatively topologically principal with respect to a family $\mathscr{H}$ of open subgroupoids then a representation of $C^{*}_{r}(G)$ is faithful if and only if it is faithful on $C^{*}_{r}(H)$ for each $H\in\mathscr{H}$. Our notion of relative topological principality is a generalization of effectiveness or topological principality for étale groupoids, but our main result also captures several cases of the ideal intersection property that had been previously studied outside of the context of effective groupoids, including results from [brown2015cartan], [LS], and [Starling].
@@
Acknowledgements: The results of this paper form part of the the second named author’s MSc Thesis [GG-MSCthesis2023] written under the supervision of the first and third named authors. We would like to thank Jack Spielberg for several helpful comments.
@@ 3. Ideal property and uniqueness theorems
Assisted by this intuition, Exel, Laca, and Quigg [ExelLacaQuigg] generalize the result of Kawamura and Tomiyama on a different axis than that of Archbold and Spielberg. They show that if a discrete group $G$ has a topologically free partial action on a space $X$ then $C_{0}(X)$ has the ideal intersection property in the reduced crossed product $C_{0}(X)\rtimes_{r}G$. Under a countability assumption this result generalizes nicely to C*-algebras of Hausdorff étale groupoids [exelNonHausdorffEtaleGroupoids2011]. For general étale groupoids the concept of topological freeness splits into several non-equivalent notions; these agree if the groupoid is second-countable and Hausdorff. Perhaps the most commonly used condition in the second-countable Hausdorff setting is the property of being topologically principal or effective which says that the interior of the isotropy is just the unit space.
@@
  • $G$ is the action groupoid of a discrete group acting (or partially acting) topologically freely on a locally compact Hausdorff space and $H$ is the unit space [ExelLacaQuigg] [KawamuraTomiyama].
  • $G$ is a topologically principal second-countable Hausdorff étale groupoid and $H$ is the unit space [exelNonHausdorffEtaleGroupoids2011].
  • $G$ is any second-countable Hausdorff étale groupoid and $H=\operatorname{Iso}(G)^{\circ}$ [brown2015cartan].
  • G is the partial transformation groupoid associated to a group embeddable monoid $P$ and $H$ is the transformation groupoid obtained by restricting the action to the group of invertible elements of $P$ [LS].
Motivated by this we aim to find a condition on open subgroupoids of a groupoid $G$ that guarantees that when taken together, their reduced C*-algebras detect ideals in the reduced C*-algebra of $G$.
@@ 4. Relative topological principality
With an eye on exploiting the orbit structure in the unit space in order to simplify ideal detection, we now introduce our main concept.
@@
It will be convenient to have a reciprocal terminology; when $G$ is topologically free relative to a family $\mathscr{H}$ of open subgroupoids we say that $\mathscr{H}$ constrains the isotropy of $G$. If the family is a singleton =$\mathscr{H}=\{H\}$, as will often be the case, we will simply say that $G$ is topologically principal relative to $H$, or that $H$ constrains the isotropy of $G$, and write $X_{H}:=X_{\{H\}}$.
@@
Let $G$ be a second-countable Hausdorff étale groupoid and suppose that $G$ is topologically principal relative to a family $\mathscr{H}$ of open subgroupoids . Then a representation $\rho$ of $C^{*}_{r}(G)$ is faithful if and only if its restriction to $C^{*}_{r}(H)\subseteq C^{*}_{r}(G)$ is faithful for each $H\in\mathscr{H}$.
To prove Theorem 4.6 we will need to use isomorphisms of C*-algebras of subgroupoids that are inner in an appropriate sense. It is natural that in a groupoid each element $\gamma\in G$ will act as a partial symmetry, but strictly speaking, one can only conjugate $\alpha\in G$ by $\gamma$ if $\textbf{s}(\gamma)=\textbf{t}(\alpha)$ and $\textbf{s}(\alpha)=\textbf{t}(\gamma^{-1})$, in which case $\textbf{s}(\alpha)=\textbf{t}(\alpha)$, so $\alpha$ must be in the isotropy bundle of $G$. This will not be sufficient for our purposes so we broaden our understanding of what it means for $G$ to act on itself by conjugation by allowing an action of $\operatorname{Bis}(G)$ on the category Sub$(G)$ of subgroupoids of $G$. If $G$ is a group then $\operatorname{Bis}(G)$ may essentially be identified with $G$ and one may recover the inner automorphisms of $G$ from this action.
We formalize this notion in the following definition.
@@
Suppose $\rho$ is faithful on $B$ and let $\tilde{\rho}$ denote the unique weak-* continuous extension to $A^{**}$. Let $c\in C$ be arbitrary and let $b\in B$ satisfy $c=vbv^{*}=\phi(b)$, where $v\in A^{**}$ is the partial isometry implementing $\phi$. Note that $\left\lVert b\right\rVert=\left\lVert c\right\rVert$ because $\phi$ is an isomorphism. Since $\rho$ is a ***-homomorphism $\left\lVert\rho(c)\right\rVert\leq\left\lVert c\right\rVert$. As for the reverse inequality we have
$\left\lVert c\right\rVert=\left\lVert b\right\rVert=\left\lVert\rho(b)\right\rVert=\left\lVert\rho(v^{*}vbv^{*}v)\right\rVert=\left\lVert\tilde{\rho}(v^{*})\rho(vbv^{*})\tilde{\rho}(v)\right\rVert\leq\left\lVert\rho(vbv^{*})\right\rVert=\left\lVert\rho(c)\right\rVert;$
@@
$(f*g)(x)=\begin{cases}f(y)g(y^{-1}x)&\text{if there exists a (necessarily unique) }y\in\Gamma\text{ such that }y^{-1}x\in V,\\ 0&\text{otherwise.}\end{cases}$
@@
proving part (3). To see that the map $\operatorname{Bis}(\Gamma)\rightarrow S$, $\Gamma\mapsto v_{\Gamma}$ is a semigroup homomorphism observe that
@@
The following two lemmas generalize results from [brown2015cartan] to a collection of subalgebras and subgroupoids.
Lemma 4.14 (cf. [brown2015cartan] Theorem 3.2. ).
@@
$\langle\pi_{\tilde{\varphi}}(x^{*}x)\pi_{\tilde{\varphi}}(a)\xi_{\tilde{\varphi}},\pi_{\tilde{\varphi}}(a)\xi_{\tilde{\varphi}}\rangle=\langle\pi_{\tilde{\varphi}}(a^{*}x^{*}xa)\xi_{\tilde{\varphi}},\xi_{\tilde{\varphi}}\rangle=\tilde{\varphi}(a^{*}x^{*}xa)=\varphi(\gamma(a^{*}x^{*}xa))=0$
@@
Lemma 4.15 (cf. [brown2015cartan] Lemma 3.3.).
Let $G$ be a second-countable Hausdorff étale groupoid that is topologically principal relative to a family $\mathscr{H}$ of open subgroupoids. Let $X_{\mathscr{H}}$ be the set defined in (4.2) and suppose that $u\in X_{\mathscr{H}}$ and $\Gamma\in\operatorname{Bis}(G)$ satisfy $u\in\Gamma^{-1}\Gamma$ and $\Gamma G^{u}_{u}\Gamma^{-1}\subseteq H$ for some $H\in\mathscr{H}$. Then for each $f\in C_{c}(G)$ there exists $b\in C_{c}((\Gamma^{-1}H\Gamma)^{(0)})^{+}$ such that $\lVert b\rVert=b(u)=1$ and $bfb\in C_{c}(\Gamma^{-1}H\Gamma)$.
@@
[…]
@@
We are now ready for the proof of our main result.
@@
$\left\lVert\vartheta_{u}(c)\right\rVert=\sup\big{\{}|\psi(\vartheta_{u}(c))|\;|\;\psi\text{ is a state on }C^{*}_{r}(G^{u}_{u})\big{\}}\leq\sup\big{\{}|\varphi(c)|\;|\;\varphi\in S_{u,\Gamma,H}\big{\}}\leq\epsilon/2.$
@@ 5. C*-algebras of Small Categories
([LiGarsideCats, Definition 2.2]) The left reduced C*-algebra of the LCSC $\mathcal{C}$ is
@@
On our way towards Spielberg’s groupoid model for $C^{*}_{r}(\mathcal{C})$ we will use the inverse semigroup approach to $C^{*}_{r}(\mathcal{C})$ described in [LiGarsideCats] because it allows us to make easy comparison with semigroup C*-algebras. Thus we turn our attention to a natural inverse semigroup arising from a LCSC.
@@
[…]
@@
One easily verifies that each $\chi_{c}\in\Omega$. Recall that a basic open subset of the spectrum $\widehat{E}$ of the semilattice $E$ is of the form $\widehat{E}(e,F):=\{\chi\;|\;\chi(e)=1,\chi(f)=0\text{ for all }f\in F\}$. Density follows immediately from the observation that if $\widehat{E}(e,F)\neq\emptyset$ then $e\setminus\cup F\neq\emptyset$, in which case $\chi_{c}\in\widehat{E}(e,F)$ for each $c\in e\setminus\cup F$. ∎
@@
$(s\cdot\chi)\left(\bigvee_{i=1}^{n}e_{i}\right)=\chi\left(s^{-1}\bigvee_{i=1}^{n}e_{i}s\right)=\chi\left(\bigvee_{i=1}^{n}s^{-1}e_{i}s\right)=\bigvee_{i=1}^{n}\chi(s^{-1}e_{i}s)=\bigvee_{i=1}^{n}(s\cdot\chi)(e_{i})$
so $\chi\in\Omega$ as desired. ∎
@@
(cf. [LiGarsideCats, Remark 2.16]) Let ¯$\overline{\mathcal{I}}_{l}*\Omega=\{(s,\chi)\in\overline{\mathcal{I}}_{l}\times\Omega\;|\;\chi(s^{-1}s)=1\}$, and define an equivalence relation on ¯$\overline{\mathcal{I}}_{l}*\Omega$ by setting
$(s,\chi)\sim(t,\psi)\iff\begin{cases}\chi=\psi\text{ and}\\ \text{there is an idempotent }e\in\overline{\mathcal{I}}_{l}\text{ such that }\chi(e)=1\text{ and }se=te.\end{cases}$
@@
For each $c\in\mathcal{C}$, the isotropy at $\chi_{c}$ is $\mathscr{G}^{\chi_{c}}_{\chi_{c}}=\{[cuc^{-1},\chi_{c}]\;|\;u\in\textbf{s}(c)\mathcal{C}^{*}\textbf{s}(c)\}$.
@@
$\mathscr{G}_{\chi_{e}}^{\chi_{e}}=\{[u,\chi_{e}]\;|\;u\in\mathcal{I}_{l},u^{-1}u=e=uu^{-1}\}=\{[u,\chi_{e}]\;|\;u\in\mathcal{C}^{*},u^{-1}u=e=uu^{-1}\},$
@@
Given $c\in\mathcal{C}$ the subset $[c,\Omega(c^{-1}c)]$ of $\mathscr{G}$ is an open bisection with $\textbf{s}([c,\Omega(c^{-1}c)])=\Omega(c^{-1}c)$ and $\textbf{t}([c,\Omega(c^{-1}c)])=\Omega(cc^{-1})$. Let $\Gamma:=[c,\Omega(c^{-1}c)]^{-1}$. Then $\Gamma\mathscr{G}^{\chi_{c}}_{\chi_{c}}\Gamma^{-1}=\mathscr{G}^{\chi_{\textbf{s}(c)}}_{\chi_{\textbf{s}(c)}}\subseteq\mathcal{C}^{*}\ltimes\Omega$; that is, the isotropy at each element of $X:=\{\chi_{c}\;|\;c\in\mathcal{C}\}$ may be conjugated into $\mathcal{C}^{*}\ltimes\Omega$ by an open bisection. Since $X$ is dense by Lemma 5.8, $\mathscr{G}$ is topologically principal relative to $\mathcal{C}^{*}\ltimes\Omega$. ∎
@@
Let $\mathcal{C}$ be a left cancellative small category such that $\mathscr{G}:=\overline{\mathcal{I}_{l}}\ltimes\Omega$ is second-countable and Hausdorff. A representation of $C_{r}^{*}(\mathcal{C})\cong C^{*}_{r}(\overline{\mathcal{I}_{l}}\ltimes\Omega)$ is faithful if and only if its restriction to the C*-subalgebra $C^{*}_{r}(\mathcal{C}^{*}\ltimes\Omega)$ is faithful.
@@
Let $\mathcal{C}$ be a left cancellative small category such that $\mathscr{G}:=\overline{\mathcal{I}_{l}}\ltimes\Omega$ is second-countable and Hausdorff. If the canonical map $C^{*}(\overline{\mathcal{I}_{l}}\ltimes\Omega)\rightarrow C^{*}_{r}(\overline{\mathcal{I}_{l}}\ltimes\Omega)$ is an isomorphism, e.g. if $\overline{\mathcal{I}_{l}}\ltimes\Omega$ is amenable, then the C*-algebra generated by a representation of $C^{*}(\overline{\mathcal{I}_{l}}\ltimes\Omega)$ that is faithful on $C^{*}(\mathcal{C}^{*}\ltimes\Omega)$ is canonically unique.

### v3

@@ Abstract.
1 Supported by NSERC Discovery Grant RGPIN-2021-02459
2 Supported by NSERC Discovery Grant RGPIN-2017-04052
@@ 1. Introduction
One often wants a method for determining the injectivity of *-homorphisms with domain a fixed C*-algebra $A$. In some situations there may be a C*-subalgebra $B$ of $A$ such that any *-homomorphism is injective on $A$ if and only if it is injective on $B$; this is particularly useful if the subalgebra $B$ is more easily understood than the whole algebra $A$. Equivalently, we say that a C*-subalgebra $B$ of $A$ detects ideals in $A$, or has the ideal intersection property in $A$, if the only (closed 2-sided) ideal $J$ of $A$ such that $B\cap J=\{0\}$ is the zero ideal. If $B$ detects ideals in $A$, then we can verify injectivity of a *-homomorphism on $A$ by checking that it is injective on $B$. There are many cases known where certain types of C*-algebras have important subalgebras with the ideal intersection property, e.g. [KawamuraTomiyama, ArchboldSpielberg, ExelLacaQuigg, exelNonHausdorffEtaleGroupoids2011, LS, brown2015cartan, Starling, kennedyIdealIntersectionProperty2021] (see Section 3 below).
In this paper we consider the ideal intersection property for subalgebras and, more generally, families of subalgebras of $C^{*}_{r}(G)$, where $G$ is a Hausdorff étale groupoid. We introduce the notion of a groupoid being relatively topologically principal with respect to a family of open subgroupoids. Our main result is that if $G$ is relatively topologically principal with respect to a family $\mathscr{H}$ of open subgroupoids then a representation of $C^{*}_{r}(G)$ is faithful if and only if it is faithful on $C^{*}_{r}(H)$ for each $H\in\mathscr{H}$. Our notion of relative topological principality is a generalization of effectiveness or topological principality for étale groupoids, but our main result also captures several cases of the ideal intersection property that had been previously studied outside of the context of effective groupoids, including results from [brown2015cartan], [LS], and [Starling].
@@
Acknowledgements: The results in this paper form part of the the second named author’s MSc Thesis [GG-MSCthesis2023] written under the supervision of the first and third named authors. We would like to thank Jack Spielberg for several helpful comments on the thesis. We would also like to thank Sergey Neshveyev and Bartosz Kwaśniewski for independently noticing that the second countability assumption made in the first posted version of this paper was not needed for our results, and in fact not used in our proofs; consequently this assumption has been removed. In addition, we also thank Kwaśniewski for telling us how to use [kwa-mey] to remove second countability from the second example at the end of Section 3, and Neshveyev for pointing out that the theory of isotropy fibers of ideals recently developed in [Chr-Nes-Isotropy-2023] is particularly well-suited to prove results such as those obtained here, including our Theorem 4.6.
We would like to thank the anonymous referee for several helpful comments and suggestions.
@@ 3. Ideal property and uniqueness theorems
Assisted by this intuition, Exel, Laca, and Quigg [ExelLacaQuigg] generalize the result of Kawamura and Tomiyama on a different axis than that of Archbold and Spielberg. They show that if a discrete group $G$ has a topologically free partial action on a space $X$ then $C_{0}(X)$ has the ideal intersection property in the reduced product $C_{0}(X)\rtimes_{r}G$. Under a second countability assumption this result has been generalized nicely to C*-algebras of Hausdorff étale groupoids in [exelNonHausdorffEtaleGroupoids2011]. For general étale groupoids the concept of topological freeness splits into several non-equivalent notions; these agree if the groupoid is second-countable and Hausdorff. Perhaps the most commonly used condition in the second-countable Hausdorff setting is the property of being topologically principal or effective which says that the interior of the isotropy is just the unit space.
@@
  • $G$ is the action groupoid of a discrete group acting (or partially acting) topologically freely on a locally compact Hausdorff space and $H$ is the unit space, see [ExelLacaQuigg, KawamuraTomiyama].
  • $G$ is a topologically principal Hausdorff étale groupoid and $H$ is the unit space, see [exelNonHausdorffEtaleGroupoids2011] for second countable $G$, and [kwa-mey, Theorem 7.29] for general $G$.
  • $G$ is any second-countable Hausdorff étale groupoid and $H=\operatorname{Iso}{(G)^{o}}$, [brown2015cartan, Theorem 3.1]. See [Starling, Theorem 2.1] for a generalization to more general open subgroupoids of $\operatorname{Iso}{(G)^{o}}$, and [Chr-Nes-Isotropy-2023, Theorem 4.8] for a generalization that includes non-Hausdorff groupoids that admit a countable cover of open bisections.
  • $G$ is the partial transformation groupoid associated to a group embeddable monoid $P$ and $H$ is the transformation groupoid obtained by restricting the action to the group of invertible elements of $P$, see [LS, Theorem 5.1].
@@ 4. Relative topological principality
With an eye on exploiting the orbit structure in the unit space of a groupoid in order to simplify ideal detection, we now introduce our main concept.
@@
It will be convenient to have a reciprocal terminology; when $G$ is topologically principal relative to a family $\mathscr{H}$ of open subgroupoids we say that $\mathscr{H}$ constrains the isotropy of $G$. If the family is a singleton =$\mathscr{H}=\{H\}$, as will often be the case, we will simply say that $G$ is topologically principal relative to $H$, or that $H$ constrains the isotropy of $G$, and write $X_{H}:=X_{\{H\}}$.
@@
Let $G$ be a Hausdorff étale groupoid and suppose that $G$ is topologically principal relative to a family $\mathscr{H}$ of open subgroupoids . Then a representation $\rho$ of $C^{*}_{r}(G)$ is faithful if and only if its restriction to $C^{*}_{r}(H)\subseteq C^{*}_{r}(G)$ is faithful for each $H\in\mathscr{H}$.
To prove Theorem 4.6 we will need to use isomorphisms of C*-algebras of subgroupoids that are inner in an appropriate sense. It is natural that in a groupoid each element $\gamma\in G$ will act as a partial symmetry, but strictly speaking, one can only conjugate $\alpha\in G$ by $\gamma$ if $\textbf{s}(\gamma)=\textbf{t}(\alpha)$ and $\textbf{s}(\alpha)=\textbf{t}(\gamma^{-1})$, in which case $\textbf{s}(\alpha)=\textbf{t}(\alpha)$, so $\alpha$ must be in the isotropy bundle of $G$. This will not be sufficient for our purposes so we broaden our understanding of what it means for $G$ to act on itself by conjugation by allowing an action of $\operatorname{Bis}(G)$ on the category Sub$(G)$ of subgroupoids of $G$. If $G$ is a group then $\operatorname{Bis}(G)$ may essentially be identified with $G$ and one may recover the inner automorphisms of $G$ from this action. We formalize this notion in the following definition.
@@
Suppose $\rho$ is faithful on $B$ and let $\tilde{\rho}$ denote the unique weak-* continuous extension to $A^{**}$. Let $c\in C$ be arbitrary and let $b\in B$ satisfy $c=vbv^{*}=\phi(b)$, where $v\in A^{**}$ is the partial isometry implementing $\phi$. Note that $\left\lVert b\right\rVert=\left\lVert c\right\rVert$ because $\phi$ is an isomorphism. Since $\rho$ is a *-homomorphism $\left\lVert\rho(c)\right\rVert\leq\left\lVert c\right\rVert$. As for the reverse inequality we have
$\left\lVert c\right\rVert=\left\lVert b\right\rVert=\left\lVert\rho(b)\right\rVert=\left\lVert\rho(v^{*}vbv^{*}v)\right\rVert=\left\lVert\tilde{\rho}(v^{*})\rho(vbv^{*})\tilde{\rho}(v)\right\rVert\\ \leq\left\lVert\rho(vbv^{*})\right\rVert=\left\lVert\rho(c)\right\rVert;$
@@
$(f*g)(x)=\begin{cases}f(y)g(y^{-1}x)&\text{if $\exists y\in\Gamma$ (necessarily unique) such that }y^{-1}x\in V,\\ 0&\text{otherwise.}\end{cases}$
@@
proving part (3). To see that the map $\operatorname{Bis}(\Gamma)\rightarrow S$, $\Gamma\mapsto v_{\Gamma}$ is multiplicative and involutive, observe that
@@
The following two lemmas generalize results from [brown2015cartan] to a collection of subalgebras and subgroupoids; we are indebted to Kwaśniewski for pointing out that second countability is not needed for part (b) of [brown2015cartan, Lemma 3.3].
Lemma 4.14 (cf. [brown2015cartan] Theorem 3.2).
@@
$\langle\pi_{\tilde{\varphi}}(x^{*}x)\pi_{\tilde{\varphi}}(a)\xi_{\tilde{\varphi}},\pi_{\tilde{\varphi}}(a)\xi_{\tilde{\varphi}}\rangle=\langle\pi_{\tilde{\varphi}}(a^{*}x^{*}xa)\xi_{\tilde{\varphi}},\xi_{\tilde{\varphi}}\rangle\\ =\tilde{\varphi}(a^{*}x^{*}xa)=\varphi(\gamma(a^{*}x^{*}xa))=0$
@@
Lemma 4.15 (cf. [brown2015cartan] Lemma 3.3).
Let $G$ be a Hausdorff étale groupoid that is topologically principal relative to a family $\mathscr{H}$ of open subgroupoids. Let $X_{\mathscr{H}}$ be the set defined in (4.2) and suppose that $u\in X_{\mathscr{H}}$ and $\Gamma\in\operatorname{Bis}(G)$ satisfy $u\in\Gamma^{-1}\Gamma$ and $\Gamma G^{u}_{u}\Gamma^{-1}\subseteq H$ for some $H\in\mathscr{H}$. Then for each $f\in C_{c}(G)$ there exists $b\in C_{c}((\Gamma^{-1}H\Gamma)^{(0)})^{+}$ such that $\lVert b\rVert=b(u)=1$ and $bfb\in C_{c}(\Gamma^{-1}H\Gamma)$.
@@
${{C^{*}_{r}(G)}}$${{C_{0}(G^{(0)})}}$${{C^{*}_{r}(G^{x}_{x})}}$${{\mathbb{C}}}$$\scriptstyle{\tau}$$\scriptstyle{\vartheta_{x}}$$\scriptstyle{\Phi}$$\scriptstyle{\operatorname{ev}_{x}}$
@@
We are now ready for the proof of our main result. The approach we take is modelled on the proof of Theorem 3.1 in [brown2015cartan].
@@
$\left\lVert\vartheta_{u}(c)\right\rVert$ $=\sup\big{\{}|\psi(\vartheta_{u}(c))|\;|\;\psi\text{ is a state on }C^{*}_{r}(G^{u}_{u})\big{\}}$
$\leq\sup\big{\{}|\varphi(c)|\;|\;\varphi\in S_{u,\Gamma,H}\big{\}}$
$\leq\epsilon/2.$
@@
Remark 4.18.
Before we launch into applications of the main result in the following two sections, we would like to point out that, as indicated after the definition, it is often the case that a singleton $\mathscr{H}$ suffices to constrain the isotropy in a useful way. Of course, in a trivial way this is always true because it is possible to choose =$\mathscr{H}=\{\mathscr{G}\}$. The point is that by making the open subgroupoids as small as possible, maybe at the cost of taking more of them, one gets a useful criterion. As an example, consider a direct sum of full matrix algebras, which is the groupoid C*-algebra of a disjoint union of groupoids, in which a constraining family is obtained by including a trivial groupoid (consisting of a single point) in $\mathscr{H}$ for each summand.
@@ 5. C*-algebras of Small Categories
The left reduced C*-algebra of the LCSC $\mathcal{C}$ is
@@
On our way towards Spielberg’s groupoid model for $C^{*}_{r}(\mathcal{C})$ we will use the inverse semigroup approach to $C^{*}_{r}(\mathcal{C})$ described in [PO2020], see also [LiGarsideCats], because it allows us to make easy comparison with semigroup C*-algebras. Thus we turn our attention to a natural inverse semigroup arising from a LCSC.
@@
We note that when 00 is in $\mathcal{I}_{l}$ the element $T_{0}$ is the zero element of the C*-algebra.
Definition 5.6.
Let $E$ be a (meet) semilattice viewed as a semigroup of idempotents and let $\{0,1\}$ be the two element boolean algebra. A semicharacter $\chi$ is a nonzero multiplicative map $\chi:E\rightarrow\{0,1\}$. The space $\widehat{E}$ is the collection of semicharacters on $E$ with the topology of pointwise convergence.
@@
${{\operatorname{Spec(C^{*}(E))}}}$${{\widehat{E}}}$${{\operatorname{Spec(D_{r})}}}$${\Omega}$$\scriptstyle{\chi\mapsto\chi|_{E}}$$\scriptstyle{\chi\mapsto\chi|_{E}}$
@@
One easily verifies that each $\chi_{c}\in\Omega$. Recall that a basic open subset of the spectrum $\widehat{E}$ of the semilattice $E$ is of the form $\widehat{E}(e,F):=\{\chi\;|\;\chi(e)=1,\chi(f)=0\text{ for all }f\in F\}$. To see density, first note that $E$ is a semilattice which is concretely represented as a subset of the power set of $\mathcal{C}$, thus we can view elements of $E$ as sets. Density then follows immediately from the observation that if $\widehat{E}(e,F)\neq\emptyset$ then $e\setminus\cup_{f\in F}f\neq\emptyset$, in which case $\chi_{c}\in\widehat{E}(e,F)$ for each $c\in e\setminus\cup_{f\in F}f$. ∎
@@
$(s\cdot\chi)\left(\bigvee_{i=1}^{n}e_{i}\right)=\chi\left(s^{-1}\bigvee_{i=1}^{n}e_{i}s\right)=\chi\left(\bigvee_{i=1}^{n}s^{-1}e_{i}s\right)\\ =\bigvee_{i=1}^{n}\chi(s^{-1}e_{i}s)=\bigvee_{i=1}^{n}(s\cdot\chi)(e_{i})$
so $\chi\in\Omega$ as desired. Note that the uses of distributivity in this calculation are justified by [lawsonInverseSemigroups1998, Proposition 1.2.2] and the fact that the above joins exist in $E$. ∎
@@
( [Paterson1999, p. 140], see also [Exel2008, Definition 4.6]) Let ¯$\overline{\mathcal{I}}_{l}*\Omega=\{(s,\chi)\in\overline{\mathcal{I}}_{l}\times\Omega\;|\;\chi(s^{-1}s)=1\}$, and define an equivalence relation on ¯$\overline{\mathcal{I}}_{l}*\Omega$ by setting
$(s,\chi)\sim(t,\psi)\iff\begin{cases}\chi=\psi\text{ and}\\ se=te\text{ for some idempotent }e\in\overline{\mathcal{I}}_{l}\text{ with }\chi(e)=1.\end{cases}$
@@
For each $c\in\mathcal{C}$, the isotropy at $\chi_{c}$ is
$\mathscr{G}^{\chi_{c}}_{\chi_{c}}=\{[cuc^{-1},\chi_{c}]\;|\;u\in\textbf{s}(c)\mathcal{C}^{*}\textbf{s}(c)\}.$
@@
$\mathscr{G}_{\chi_{e}}^{\chi_{e}}$ =$=\{[u,\chi_{e}]\;|\;u\in\mathcal{I}_{l},u^{-1}u=e=uu^{-1}\}$
=$=\{[u,\chi_{e}]\;|\;u\in\mathcal{C}^{*},u^{-1}u=e=uu^{-1}\},$
@@
Given $c\in\mathcal{C}$ the subset $[c,\Omega(c^{-1}c)]$ of $\mathscr{G}$ is an open bisection with $\textbf{s}([c,\Omega(c^{-1}c)])=\Omega(c^{-1}c)$ and $\textbf{t}([c,\Omega(c^{-1}c)])=\Omega(cc^{-1})$. If we let $\Gamma:=[c,\Omega(c^{-1}c)]^{-1}$, then $\Gamma\mathscr{G}^{\chi_{c}}_{\chi_{c}}\Gamma^{-1}=\mathscr{G}^{\chi_{\textbf{s}(c)}}_{\chi_{\textbf{s}(c)}}\subseteq\mathcal{C}^{*}\ltimes\Omega$; that is, the isotropy at each element of $X:=\{\chi_{c}\;|\;c\in\mathcal{C}\}$ may be conjugated into $\mathcal{C}^{*}\ltimes\Omega$ by an open bisection. Since $X$ is dense by Lemma 5.9, $\mathscr{G}$ is topologically principal relative to $\mathcal{C}^{*}\ltimes\Omega$. ∎
@@
Let $\mathcal{C}$ be a left cancellative small category such that $\mathscr{G}:=\overline{\mathcal{I}_{l}}\ltimes\Omega$ is Hausdorff. A representation of $C_{r}^{*}(\mathcal{C})\cong C^{*}_{r}(\overline{\mathcal{I}_{l}}\ltimes\Omega)$ is faithful if and only if its restriction to the C*-subalgebra $C^{*}_{r}(\mathcal{C}^{*}\ltimes\Omega)$ is faithful.
@@
Let $\mathcal{C}$ be a left cancellative small category such that $\mathscr{G}:=\overline{\mathcal{I}_{l}}\ltimes\Omega$ is Hausdorff. If the canonical map $C^{*}(\overline{\mathcal{I}_{l}}\ltimes\Omega)\rightarrow C^{*}_{r}(\overline{\mathcal{I}_{l}}\ltimes\Omega)$ is an isomorphism, e.g. if $\overline{\mathcal{I}_{l}}\ltimes\Omega$ is amenable, then the C*-algebra generated by a representation of $C^{*}(\overline{\mathcal{I}_{l}}\ltimes\Omega)$ that is faithful on $C^{*}(\mathcal{C}^{*}\ltimes\Omega)$ is canonically unique.